OpenStudy (anonymous):

Simplify these expressions. Use only positive exponents. p^-5 1/7t^-3 (6m^-2n^3)^0

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... just the first one is almost a given one recall that \(\bf a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}} \\ \quad \\ % negative exponential denominator a^{{\color{red} n}} \implies \cfrac{1}{a^{-\color{red} n}} \qquad \qquad \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}} \\ \quad \\ % rational with exponential terms \cfrac{a^{\color{red}{ m}}}{a^{\color{purple}{ n}}}\implies \cfrac{a^{\color{red}{ m}}}{1}\cdot \cfrac{1}{a^{\color{purple}{ n}}}\implies a^{\color{red}{ m}}\cdot a^{-\color{purple}{ n}}\implies a^{{\color{red}{ m}}{-\color{purple}{ n}}}\)

OpenStudy (anonymous):

It just all looks like gibberish....

OpenStudy (jdoe0001):

hheheh w0t?

OpenStudy (jdoe0001):

are you hmm using Internet Explorer by any chance? or you can see the red numbers?

OpenStudy (anonymous):

I can see them. I just don't understand them.

OpenStudy (jdoe0001):

the "a" is any value pretty much raised to some "n" exponent if the exponent is negative, you flip it over the fraction, to get a positive exponent :)

OpenStudy (jdoe0001):

so \(\bf p^{-5}\implies \cfrac{1}{p^{\color{red}{ 5}}}\)

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{7t^{-3}}\implies \cfrac{1}{7}\cdot \cfrac{1}{t^{-3}}\implies \cfrac{1}{7}\cdot t^{\color{red}{ 3}}\implies \cfrac{t^3}{7}\)

OpenStudy (jdoe0001):

the last one is raised to the "0" exponent, and well.. anything raised to 0 (except 0 itself), is 1 :)

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

Do you think you could help me with the last one please?

zepdrix (zepdrix):

hi book c:

OpenStudy (anonymous):

hi! :)

zepdrix (zepdrix):

When you raise something to the 1st power, you leave it alone. \(\rm (stuff)^1=stuff\) When you raise something to the 2nd power, you multiply it by a copy of itself. \(\rm (stuff)^2=(stuff)(stuff)\) So when you raise something to the 0th power, you're actually dividing the only copy out of itself. \(\rm (stuff)^0=(stuff)\div(stuff)=1\)

zepdrix (zepdrix):

Anything* divided by itself, gives us 1, ya? Because it divides evenly into itself one time.

OpenStudy (anonymous):

yeah, but i'm not sure I see where this is going...

OpenStudy (anonymous):

(sorry i'm just about clueless when it comes to math.)

zepdrix (zepdrix):

The point is that,\[\large\rm (6m^{-2}n^3)^0=(6m^{-2}n^3)\div(6m^{-2}n^3)\]\[\large\rm (6m^{-2}n^3)^0=1\]you don't actually have to distribute the 0 to each item, the 0 power always give us 1 as a result :)

OpenStudy (anonymous):

So it looks long (maybe a bit complicated) and it only ends up as a 1 because of the 0th power?

zepdrix (zepdrix):

Yes :)\[\large\rm 4^0=1\]\[\large\rm (7x)^0=1\]\[\large\rm (12x^2y^3)^0=1\]\[\large\rm (75a^{-4}b^{3}c^{54}d^{12})^0=1\]

zepdrix (zepdrix):

Keep in mind that ONLY the part with the 0th power on it turns into a 1 though, Example:\[\large\rm 7\color{orangered}{(12x^2)^0}=7\cdot \color{orangered}{1}\]

OpenStudy (anonymous):

Good to know! Thank you so much! I think that because of your help (and the time you took to explain things so I understood.) I'm going to have to knowledge to pass my test next week. I really don't think I can thank you enough!

zepdrix (zepdrix):

lol aw c:

zepdrix (zepdrix):

Keep up the hard work, you'll do great. Be blessed! c: